OpenStudy (anonymous):

if the internal energy of a thermodynamic system is increased by 300 J while 75 J of expansion work is done, how much heat was transferred and in which direction, to or from the system

OpenStudy (surry99):

Hint: What is the 1st law of thermodynamics?

OpenStudy (anonymous):

energy is conserved

OpenStudy (surry99):

Agreed...what is it mathematically in terms of internal energy, heat and work?

OpenStudy (anonymous):

'triangle'=q+w?

OpenStudy (anonymous):

"triangle E"

OpenStudy (anonymous):

change in enthalpy=work + heat

OpenStudy (surry99):

so: delta U = q + w (I use U you can use E if you like)

OpenStudy (surry99):

No not enthalpy...it say s change in internal energy = heat + work

OpenStudy (anonymous):

ya, i know the equations, im asking an explanation of the energy transfer and yeah, my bad, its internal energy

OpenStudy (surry99):

In this problem which two values are you given?

OpenStudy (anonymous):

three hunna and 75, both J, the system adds 300, buts 75 J of expansion work is done i dont understand what they are saying with "75 J of expansion work"

OpenStudy (surry99):

ok... by definition: if heat flows into the system it is positive, if it flow out it is negative if the system does work on the surroundings it is negative, if the surroundings does work on the system it is positive so the 75J of expansion work means the system is doing work on the surroundings therefore by convention the sign for that is negative so: delta E = q + w 300 J = q + (-75J)

OpenStudy (surry99):

Check your sign conventions in your notes.

OpenStudy (anonymous):

so they expansion means the system is doing the work to the surroundings?

OpenStudy (anonymous):

the* not they

OpenStudy (surry99):

exactly my friend...the system is doing work on the surroundings

OpenStudy (anonymous):

ok, the "expansion" confused me

OpenStudy (anonymous):

thank you

OpenStudy (surry99):

you are welcome...don't forget to check your notes for the work sign convention you are using in your class

OpenStudy (anonymous):

k, got it